# On the cardinality of measures of maximal relative entropy for smooth skew products

### v2

@@ 4.1. Petersen-Quas-Shin construction
Diagram 2: Diagram of the maps $p_{X}^{i},\mathrm{proj}_{Z_{n}}$, $\mathrm{proj}_{X_{n}}$, $\pi$, $\pi_{n}$, $\mathrm{proj}_{X}^{i}$ and $\mathrm{proj}^{i}_{X_{n}}$ with $i\in\{1,2,3\}$. The right diagram commutes, and all maps are surjective semi-conjugacies between the shift maps on the respective spaces.
@@
Below we quote another result from [44]. In [44], the statement is proved under the assumption that $X$ is a subshift of finite type, but the same proof works assuming that $X$ is a countable-state irreducible Markov shift and $Y$ is a subshift of finite type.
Lemma 4.8 ([44, Lemma 3]).
@@
Consider a two-sided countable Markov shift $X\subset\mathcal{A}_{X}^{\mathbb{Z}}$ with shift $\sigma_{X}$, a subshift $Y$ with shift $\sigma_{Y}$, and a $1$-block factor map $\pi:X\to Y$. Fix an ergodic $\sigma_{Y}$-invariant probability $\nu$ on $Y$ and choose ergodic $\mu_{1},\mu_{2}\in\pi^{-1}\{\nu\}$. Denote by $\widehat{\mu}:=\mu_{1}\otimes_{\nu}\mu_{2}$ their relatively independent joining on $X\times X$, and by $p_{X}^{1},p_{X}^{2}:X\times X\to X$ the coordinate projections.
@@
for every $j\in\mathcal{A}_{X}$, then $\mu_{1}=\mu_{2}$.
@@
We assume the notations of Theorem 4.6 and Lemma 4.8. Assume, for contradiction, that the theorem is false. Then there exist two ergodic measures of maximal $\nu$-relative entropy, $\mu^{1}$ and $\mu^{2}$, such that
@@
Define $\mu^{3}=(\mathrm{proj}^{3}_{X})_{*}\eta$ by the construction at the beginning of Section 4.1 (we keep the same notation, omitting the index $n$). We will show that
@@
which yields a contradiction. Note that we cannot directly apply the result of Lemma 4.8 since $X$ is not a subshift of finite type.
From the discussion following the proof of Corollary 4.5,
@@
Since $\mu^{1}\otimes_{\nu}\mu^{2}$ is invariant under $\sigma_{X}\times\sigma_{X}$,
$\mu^{1}\otimes_{\nu}\mu^{2}\left(\{(u,v)\in X\times X;\,u_{-1}=v_{-1}\}\right)>0.$
By Lemma 4.8 there exists $j_{0}\in\mathcal{A}_{X}$ such that, on a $\mu^{1}\otimes_{\nu}\mu^{2}$-positive subset of $\{(u,v)\in X\times X:u_{-1}=v_{-1}\}$,
$\mathbb{E}_{\mu^{1}\otimes_{\nu}\mu^{2}}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ p_{X}^{1}\,\bigm|\,\mathcal{F}^{-}_{1}\vee\mathcal{B}_{0}\bigr)\neq\mathbb{E}_{\mu^{1}\otimes_{\nu}\mu^{2}}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ p_{X}^{2}\,\bigm|\,\mathcal{F}^{-}_{2}\vee\mathcal{B}_{0}\bigr).$
@@
$\mathbb{E}_{\eta}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ\mathrm{proj}_{X}^{1}\,\bigm|\,\mathcal{B}^{-}_{1}\vee\mathcal{B}_{0}\bigr)\neq\mathbb{E}_{\eta}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ\mathrm{proj}_{X}^{2}\,\bigm|\,\mathcal{B}^{-}_{2}\vee\mathcal{B}_{0}\bigr)$ (10)
on an $\eta$-positive measure set contained in $S:=\{\,z\in Z;\ (\mathrm{proj}_{X}^{1}(z))_{-1}=(\mathrm{proj}_{X}^{2}(z))_{-1}\,\}.$
@@
$\sigma\left(\bigcup_{n\geq 1}(\mathrm{proj}_{Z_{n}})^{-1}\bigl(\mathcal{B}_{i,n}^{-}\vee\mathcal{B}_{0,n}\bigr)\right)=\mathcal{B}_{i}^{-}\vee\mathcal{B}_{0}.$
@@
$\mathbb{E}_{\eta}\bigl(\mathbbm{1}_{[j_{0}]_{0}}\circ\mathrm{proj}_{X}^{i}\,\bigm|\,\mathcal{B}^{-}_{i}\vee\mathcal{B}_{0}\bigr)=\lim_{n\to\infty}\mathbb{E}_{\eta}\left(\mathbbm{1}_{[j_{0}]_{0}}\circ\mathrm{proj}_{X}^{i}\,\bigm|\,(\mathrm{proj}_{Z_{n}})^{-1}(\mathcal{B}_{i,n}^{-}\vee\mathcal{B}_{0,n})\right)$ (11)
@@
$\mathrm{proj}_{Z_{n}}(S)\subset S_{n}:=\{\,z\in Z_{n};\ (\mathrm{proj}_{X_{n}}^{1}(z))_{-1}=(\mathrm{proj}_{X_{n}}^{2}(z))_{-1}\,\}.$
@@
$\int_{S_{n}}\psi\left(\frac{A_{n}^{j_{0}}+B_{n}^{j_{0}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}})-\frac{1}{2}\psi(B_{n}^{j_{0}})\,\mathrm{d}\eta_{n}$
$=\int_{{\mathrm{proj}^{-1}_{Z_{n}}}(S_{n})}\psi\left(\frac{A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}+B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}})-\frac{1}{2}\psi(B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}})\,\mathrm{d}\eta$
$\geq\int_{S}\psi\left(\frac{A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}+B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}})-\frac{1}{2}\psi(B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}})\,\mathrm{d}\eta,$ (12)
where $A_{n}^{j},B_{n}^{j}$ are as in Theorem 4.6. Observe that
$A_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}$ $=\mathbb{E}_{\eta_{n}}\big(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\,\big|\,\mathcal{B}_{1,n}^{-}\vee\mathcal{B}_{0,n}\big)\circ\mathrm{proj}_{Z_{n}}$
$=\mathbb{E}_{\eta}\left(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X}^{1}\,\big|\,(\mathrm{proj}_{Z_{n}})^{-1}\left(\mathcal{B}_{1,n}^{-}\vee\mathcal{B}_{0,n}\right)\right)$ (13)
@@
$B_{n}^{j_{0}}\circ\mathrm{proj}_{Z_{n}}=\mathbb{E}_{\eta}\left(\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X}^{2}\,\big|\,(\mathrm{proj}_{Z_{n}})^{-1}\left(\mathcal{B}_{2,n}^{-}\vee\mathcal{B}_{0,n}\right)\right).$ (14)
@@
$\lim_{n\to\infty}\int_{S_{n}}\left[\psi\left(\frac{A_{n}^{j_{0}}+B_{n}^{j_{0}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{j_{0}})-\frac{1}{2}\psi(B_{n}^{j_{0}})\right]\mathrm{d}\eta_{n}>0.$ (15)

### v3

@@ 4.1. Petersen-Quas-Shin construction
Diagram 2: Diagram of the maps $p_{X}^{i},\mathrm{proj}_{Z_{n}}$, $\mathrm{proj}_{X_{n}}$, $\pi$, $\pi_{n}$, $\mathrm{proj}_{X}^{i}$ and $\mathrm{proj}^{i}_{X_{n}}$ with $i\in\{1,2,3\}$. The right diagram commutes, and all maps are semi-conjugacies between the shift maps on the respective spaces.
Sketch of the proof of Theorem 4.6.
We follow closely [Quas, Proof of Theorem 1]. Let $\mathcal{P}$ be the partition into 0 cylinders of $X_{n}$ and $\mathcal{P}_{i}=(\mathrm{proj}_{X_{n}}^{i})^{-1}\mathcal{P}$ for $i\in\{1,2,3\}$. For simplicity of the notation, since $n$ is fixed we write $\mathcal{B}_{1}^{-}:=\mathcal{B}_{1,n}^{-},\mathcal{B}_{2}^{-}:=\mathcal{B}_{2,n}^{-},$
$\mathcal{B}_{3}^{-}=\mathcal{B}_{3,n}^{-}=\sigma\left(\left\{z\in Z_{n};(\mathrm{proj}_{X_{n}}^{3}(z))_{-m}=j\right\};j\in\mathcal{A}_{X_{n}}\ \text{and }m\in\mathbb{N}\right)$
and $\mathcal{B}_{0,n}:=\mathcal{B}_{0}.$
First of all, for each $i\in\{1,2,3\}$ we have that (see [Quas, Equation (33)])
$h_{\mu_{n}^{i}}(\sigma_{X_{n}})=H_{\eta_{n}}\left(\mathcal{P}_{i}\mid\mathcal{B}_{i}^{-}\vee\mathcal{B}_{0}\right)+h_{\nu}(Y).$
By definition,
$H_{\eta_{n}}(\mathcal{P}_{i}\mid\mathcal{B}_{i}^{-}\vee\mathcal{B}_{0})$ $=\int_{Z_{n}}-\sum_{j\in\mathcal{A}_{X_{n}}}\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{i}\log\left(\mathbb{E}_{\eta_{n}}\left[\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{i}\mid\mathcal{B}_{i}^{-}\vee\mathcal{B}_{0}\right]\right)\mathrm{d}\eta_{n}$
$\hskip-56.9055pt=\int_{Z_{n}}-\sum_{j\in\mathcal{A}_{X_{n}}}\mathbb{E}_{\eta_{n}}\left[\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{i}\mid\mathcal{B}_{i}^{-}\vee\mathcal{B}_{0}\right]\log\left(\mathbb{E}_{\eta_{n}}\left[\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{i}\mid\mathcal{B}_{i}^{-}\vee\mathcal{B}_{0}\right]\right)\mathrm{d}\eta_{n}$
$=\int_{Z_{n}}\sum_{j\in\mathcal{A}_{X_{n}}}\psi\left(\mathbb{E}_{\eta_{n}}\left[\mathbbm{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{i}\mid\mathcal{B}_{i}^{-}\vee\mathcal{B}_{0}\right]\right)\mathrm{d}\eta_{n},$ (15)
where $\psi(t)=-t\log(t).$
Repeating verbatim the computation in [Quas, Proof of identity (35)], we obtain that for any $j\in\mathcal{A}_{X_{n}}$ the function
$g_{n}^{j}({z}):=\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{3}\mid\mathcal{B}_{1}^{-}\vee\mathcal{B}_{2}^{-}\vee\mathcal{B}_{3}^{-}\vee\mathcal{B}_{0}\big)({z}),$
where z=(u,v,r),{z}=({u},{v},{r}), satisfies
  • If $(\mathrm{proj}_{X_{n}}^{3}(z))_{-1}=u_{-1}\neq v_{-1}$, then
    $g_{n}^{j}(z)=\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\mid\mathcal{B}_{1}^{-}\vee\mathcal{B}_{0}\big)(z)=A^{j}_{n}(z).$
  • If $(\mathrm{proj}_{X_{n}}^{3}(z))_{-1}=v_{-1}\neq u_{-1}$, then
    $g_{n}^{j}(z)=\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{2}\mid\mathcal{B}_{2}^{-}\vee\mathcal{B}_{0}\big)(z)=B^{j}_{n}(z).$
  • If $u_{-1}=v_{-1}=a$, then
    $g_{n}^{j}(z)$ $=\frac{1}{2}\,\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\mid\mathcal{B}_{1}^{-}\vee\mathcal{B}_{0}\big)(z)+\frac{1}{2}\,\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{2}\mid\mathcal{B}_{2}^{-}\vee\mathcal{B}_{0}\big)(z)$
    $=\frac{1}{2}\left(A_{n}^{j}(z)+B_{n}^{j}(z)\right)$
  • If $u_{-1}=v_{-1}\neq a$, then
    $g_{n}^{j}(z)=\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\mid\mathcal{B}_{1}^{-}\vee\mathcal{B}_{0}\big)(z)\ \text{when }r_{N_{-1}^{n}(u,v)}=0$
    and
    $g_{n}^{j}(z)=\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{2}\mid\mathcal{B}_{2}^{-}\vee\mathcal{B}_{0}\big)(z)\ \text{when }r_{N_{-1}^{n}(u,v)}=1.$
From (15) we obtain that
$H_{\eta_{n}}(\mathcal{P}_{3}\mid\mathcal{B}_{3}^{-}\vee\mathcal{B}_{0})$ $\geq H_{\eta_{n}}(\mathcal{P}_{3}\mid\mathcal{B}_{1}^{-}\vee B_{2}^{-}\vee B_{3}^{-}\vee\mathcal{B}_{0})$ (16)
$=\int_{Z_{n}}\sum_{j\in\mathcal{A}_{X_{n}}}\psi\circ g_{n}^{j}(z)\,\mathrm{d}\eta_{n}(z)$ (17)
We define the sets
  • $S_{1}=\{z=(u,v,r)\in Z_{n};\,(\mathrm{proj}_{X_{n}}^{3}(z))_{-1}=u_{-1}\neq v_{-1}\}$;
  • $S_{2}=\{z=(u,v,r)\in Z_{n};\,(\mathrm{proj}_{X_{n}}^{3}(z))_{-1}=v_{-1}\neq u_{-1}\}$;
  • $S_{3}^{(a)}=\{z=(u,v,r)\in Z_{n};u_{-1}=v_{-1}=a\}$; and
  • $S_{3}^{(\neq a)}=\{z=(u,v,r)\in Z_{n};u_{-1}=v_{-1}\neq a\}$.
As in [Quas, Equations (41), (42) and (43)] we have that
$\int_{S_{1}\cup S_{2}}\psi\circ g_{n}^{j}(z)\mathrm{d}\eta_{n}(z)=$ $\frac{1}{2}\int_{S_{1}\cup S_{2}}\psi\big(\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\mid\mathcal{B}_{1}^{-}\vee\mathcal{B}_{0})\big)(z)\ \mathrm{d}\eta_{n}(z)$
$+\frac{1}{2}\int_{S_{1}\cup S_{2}}\psi\big(\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{2}\mid\mathcal{B}_{2}^{-}\vee\mathcal{B}_{0})\big)(z)\ \mathrm{d}\eta_{n}(z)$ (18)
Since $\eta_{n}=\mu_{n}^{1}\otimes_{\nu}\mu_{n}^{2}\otimes\zeta$ and $\zeta$ is $\left(\frac{1}{2},\frac{1}{2}\right)$-Bernoulli we have that
$\zeta(r_{N_{-1}^{n}(u,v)}=0)=\zeta(r_{N_{-1}^{n}(u,v)}=1)=\frac{1}{2}$
and we obtain that
$\int_{S_{3}^{(\neq a)}}\psi\circ g_{n}^{j}\,\mathrm{d}\eta_{n}=$ $\frac{1}{2}\int_{S_{3}^{(\neq a)}}\psi\circ\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\mid\mathcal{B}_{1}^{-}\vee\mathcal{B}_{0}\big)\,\mathrm{d}\eta_{n}$
$+\frac{1}{2}\int_{S_{3}^{(\neq a)}}\psi\circ\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{2}\mid\mathcal{B}_{2}^{-}\vee\mathcal{B}_{0}\big)\,\mathrm{d}\eta_{n}.$ (19)
Furthermore,
$\int_{S_{3}^{(a)}}\psi\circ g_{n}^{j}(z)\,\mathrm{d}\eta_{n}$ $=\int_{S_{3}^{(a)}}\psi\left(\frac{1}{2}\left(A_{n}^{j}+B_{n}^{j}\right)\right)\,\mathrm{d}\eta_{n}$ (20)
$=\int_{S_{3}^{(a)}}\frac{1}{2}\psi\circ A_{n}^{j}\,\mathrm{d}\eta_{n}+\int_{S_{3}^{(a)}}\frac{1}{2}\psi\circ B_{n}^{j}\mathrm{d}\eta_{n}+\int_{S_{3}^{(a)}}\Xi_{n}^{j}\,\mathrm{d}\eta_{n}.$ (21)
Combining (17), (18), (19) and (21) we obtain that
$h_{\mu_{3}^{n}}(\sigma_{X_{n}}\mid\nu)=$ $H_{\eta_{n}}(\mathcal{P}_{3}\mid\mathcal{B}_{3}^{-}\vee\mathcal{B}_{0})\geq\int_{Z_{n}}\sum_{j\in\mathcal{A}_{X_{n}}}\psi\circ g_{n}^{j}\,\mathrm{d}\eta_{n}$
$=$ $\int_{S_{1}\cup S_{3}\cup S_{3}^{(\neq a)}}\sum_{j\in\mathcal{A}_{X_{n}}}\psi\circ g_{n}^{j}\mathrm{d}\eta_{n}+\int_{S_{3}^{(a)}}\sum_{j\in\mathcal{A}_{X_{n}}}\psi\circ g_{n}^{j}\,\mathrm{d}\eta_{n}$
$\geq$ $\frac{1}{2}\int_{X_{n}}\psi\circ\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\mid\mathcal{B}_{1}^{-}\vee\mathcal{B}_{0}\big)\,\mathrm{d}\eta_{n}$
$+\frac{1}{2}\int_{X_{n}}\psi\circ\mathbb{E}_{\eta_{n}}\big(\mathbf{1}_{[j]_{0}}\circ\mathrm{proj}_{X_{n}}^{2}\mid\mathcal{B}_{2}^{-}\vee\mathcal{B}_{0}\big)\,\mathrm{d}\eta_{n}+\sum_{j\in\mathcal{A}_{X_{n}}}\int_{S_{3}^{(a)}}\Xi_{n}^{j}\,\mathrm{d}\eta_{n}$
$\geq\frac{h_{\mu_{1}^{n}}(\sigma_{X_{n}}\mid\nu)+h_{\mu_{2}^{n}}(\sigma_{X_{n}}\mid\nu)}{2}+\int_{S_{3}^{(a)}}\Xi_{n}^{j_{*}}\,\mathrm{d}\eta_{n}.$
@@
Below we quote another result from [Quas]. In [Quas], the statement is proved under the assumption that $X$ is a subshift of finite type, but the same proof works assuming that $X$ is a countable-state irreducible Markov shift and $Y$ is a subshift of finite type.
Lemma 4.8 ([Quas, Lemma 3]).
@@
Consider a two-sided countable Markov shift $X\subset\mathcal{A}_{X}^{\mathbb{Z}}$ with shift $\sigma_{X}$, a subshift $Y$ with shift $\sigma_{Y}$, and a $1$-block map $\pi:X\to Y$ such that $\nu(\pi(X))=1$ and $\mathrm{supp}\,(\nu)=Y$. Fix an ergodic $\sigma_{Y}$-invariant probability $\nu$ on $Y$ and choose ergodic $\mu_{1},\mu_{2}\in\pi^{-1}\{\nu\}$. Denote by $\widehat{\mu}:=\mu_{1}\otimes_{\nu}\mu_{2}$ their relatively independent joining on $X\times X$, and by $p_{X}^{1},p_{X}^{2}:X\times X\to X$ the coordinate projections.
@@
for every $j\in\mathcal{A}_{X}$, then $\mu_{1}=\mu_{2}$. In particular, there exists $a,b\in\mathcal{A}_{X}$ such that
$\widehat{\mu}\left\{({u},{v})\in X\times X\,\left|\begin{aligned} &u_{-1}=v_{-1}=a,\text{and }\\ &\mathbb{E}_{\widehat{\mu}}\bigl(\mathbbm{1}_{[b]_{0}}\circ\mathrm{proj}_{X}^{1}\,\bigm|\,\mathcal{F}^{-}_{1}\vee\mathcal{F}_{0}\bigr)\neq\mathbb{E}_{\widehat{\mu}}\bigl(\mathbbm{1}_{[b]_{0}}\circ\mathrm{proj}_{X}^{2}\bigm|\,\mathcal{F}^{-}_{2}\vee\mathcal{F}_{0}\bigr)\end{aligned}\right.\right\}>0.$
@@
Let $a,b\in\mathcal{A}_{X}$ be as given at the end of Lemma 4.8, and let $n\in\mathbb{N}$ be large enough such that $a]_{0},[b]_{0}\in\alpha_{n}$; in this way we have that $a,b\in\mathcal{A}_{X_{n}}$. We assume the notations of Theorem 4.6 and Lemma 4.8. Suppose, for contradiction, that the theorem is false. Then there exist two ergodic measures of maximal $\nu$-relative entropy, $\mu^{1}$ and $\mu^{2}$, such that
@@
Let $\eta:=\mu^{1}\otimes_{\nu}\mu^{2}\otimes\zeta$, recalling that $\zeta$ is the $\mathrm{Bernoulli}(\frac{1}{2},\frac{1}{2})$ probability measure on $\{0,1\}^{\mathbb{Z}}$. Define $\mu_{3}=(\mathrm{proj}^{3}_{X})_{*}\eta$, where for $k\in\mathbb{Z}$ we set
$N_{k}(u,v):=\sup\{m<k;\,u_{m}=v_{m}=a\}\ \text{with}\ \sup\varnothing=-\infty.$
Recalling the definition of $r_{-\infty}$ in (12) we define
$\bigl(\mathrm{proj}^{3}_{X}(u,v,r)\bigr)_{k}:=\begin{cases}u_{k},&\text{if }r_{N_{k}(u,v)}=0,\\ v_{k},&\text{if }r_{N_{k}(u,v)}=1.\end{cases}$
Observe that $\mathrm{proj}_{X}^{3}$ is the non-truncated analogue of the $\mathrm{proj}_{X_{n}}^{3}$ function in Section 4.1, obtained by replacing $X_{n}$ with $X$. We will show that
@@
which yields a contradiction. Note that we cannot directly apply the result of Lemma 4.8 since $(X,\sigma_{X})$ is not necessarily a subshift of finite type.
We define $\mu^{i}_{n}=(\mathrm{proj}_{X_{n}})_{*}\mu^{i}$ for $i\in\{1,2\}$ and we recall that $\eta_{n}=\mu_{1}^{n}\otimes_{\nu}\mu_{2}^{n}\otimes\zeta$. Let $\mu_{n}^{3}$ be the $\sigma_{X_{n}}$-invariant probability measure constructed in Section 4.1. Observe that by our choice of $n\in\mathbb{N}$ the symbol $a\in\mathcal{A}_{X}$ also lies in the alphabet of its truncation $\mathcal{A}_{X_{n}}$.
We claim that $(\mathrm{proj}_{X_{n}})_{*}\mu^{3}=\mu_{n}^{3},$ which is equivalent (from the definitions of $\mu^{3}$ and $\mu_{n}^{3}$) to showing that
$(\mathrm{proj}_{X_{n}})_{*}(\mathrm{proj}_{X}^{3})_{*}\left(\mu^{1}\otimes_{\nu}\mu^{2}\otimes\zeta\right)$ $=(\mathrm{proj}_{X_{n}}^{3})_{*}\left[(\mathrm{proj}_{X_{n}})_{*}\mu^{1}\otimes_{\nu}(\mathrm{proj}_{X_{n}})_{*}(\mu^{2})\otimes\zeta\right]$ (22)
$=(\mathrm{proj}_{X_{n}}^{3})_{*}\left(\mu^{1}_{n}\otimes_{\nu}\mu^{2}_{n}\otimes\zeta\right)$ (23)
To show (23), observe from (11) and the fact that $a\in\mathcal{A}_{X_{n}}$ we have that
$N_{k}(u,v)$ $=\sup\left\{m<k;u_{m}=v_{m}=a\right\}=\sup\left\{m<k;\mathrm{proj}_{X_{n}}(u)_{m}=\mathrm{proj}_{X_{n}}(v)_{m}=a\right\}$
$=N_{k}^{n}(\mathrm{proj}_{X_{n}}(u),\mathrm{proj}_{X_{n}}(v))$ (24)
From (24), we obtain that for every $k\in\mathbb{Z}$
$\left(\mathrm{proj}_{X_{n}}\circ\mathrm{proj}_{X}^{3}(u,v,r)\right)_{k}$ $=\begin{cases}\mathrm{proj}_{X_{n}}(u)_{k},&\text{if }r_{N_{k}(u,v)}=0,\\[2.0pt] \mathrm{proj}_{X_{n}}(v)_{k},&\text{if }r_{N_{k}(u,v)}=1.\end{cases}$
$=\begin{cases}\mathrm{proj}_{X_{n}}(u)_{k},&\text{if }r_{N_{k}^{n}(\mathrm{proj}_{X_{n}}(u),\mathrm{proj}_{X_{n}}(v))}=0,\\[2.0pt] \mathrm{proj}_{X_{n}}(v)_{k},&\text{if }r_{N_{k}^{n}(\mathrm{proj}_{X_{n}}(u),\mathrm{proj}_{X_{n}}(v))}=1.\end{cases}$
$=\mathrm{proj}_{X_{n}}^{3}(\mathrm{proj}_{X_{n}}(u),\mathrm{proj}_{X_{n}}(v),r)$
Which implies (23) and therefore that $\mathrm{proj}_{X_{n}}\mu^{3}=\mu_{n}^{3}$
In this way, we have that
@@
By Lemma 4.8 there exists $b\in\mathcal{A}_{X}$ (which also lies in $\mathcal{A}_{X_{n}}$ for $n$ large enough) such that, on a $\mu^{1}\otimes_{\nu}\mu^{2}$-positive subset of $\{(u,v)\in X\times X:u_{-1}=v_{-1}=a\}$,
$\mathbb{E}_{\mu^{1}\otimes_{\nu}\mu^{2}}\bigl(\mathbbm{1}_{[b]_{0}}\circ p_{X}^{1}\,\bigm|\,\mathcal{F}^{-}_{1}\vee\mathcal{B}_{0}\bigr)\neq\mathbb{E}_{\mu^{1}\otimes_{\nu}\mu^{2}}\bigl(\mathbbm{1}_{[b]_{0}}\circ p_{X}^{2}\,\bigm|\,\mathcal{F}^{-}_{2}\vee\mathcal{B}_{0}\bigr).$
@@
$\mathbb{E}_{\eta}\bigl(\mathbbm{1}_{[b]_{0}}\circ\mathrm{proj}_{X}^{1}\,\bigm|\,\mathcal{B}^{-}_{1}\vee\mathcal{B}_{0}\bigr)\neq\mathbb{E}_{\eta}\bigl(\mathbbm{1}_{[b]_{0}}\circ\mathrm{proj}_{X}^{2}\,\bigm|\,\mathcal{B}^{-}_{2}\vee\mathcal{B}_{0}\bigr)$ (25)
on an $\eta$-positive measure set contained in
$S^{(a)}:=\{\,z\in Z;\ (\mathrm{proj}_{X}^{1}(z))_{-1}=(\mathrm{proj}_{X}^{2}(z))_{-1}=a\}.$
@@
$\sigma\left(\bigvee_{n\geq 1}(\mathrm{proj}_{Z_{n}})^{-1}\bigl(\mathcal{B}_{i,n}^{-}\vee\mathcal{B}_{0,n}\bigr)\right)=\mathcal{B}_{i}^{-}\vee\mathcal{B}_{0}.$
@@
$\mathbb{E}_{\eta}\bigl(\mathbbm{1}_{[b]_{0}}\circ\mathrm{proj}_{X}^{i}\,\bigm|\,\mathcal{B}^{-}_{i}\vee\mathcal{B}_{0}\bigr)=\lim_{n\to\infty}\mathbb{E}_{\eta}\left(\mathbbm{1}_{[b]_{0}}\circ\mathrm{proj}_{X}^{i}\,\bigm|\,(\mathrm{proj}_{Z_{n}})^{-1}(\mathcal{B}_{i,n}^{-}\vee\mathcal{B}_{0,n})\right)$ (26)
@@
$\mathrm{proj}_{Z_{n}}(S^{(a)})\subset S_{n}^{(a)}:=\{\,z\in Z_{n};\ (\mathrm{proj}_{X_{n}}^{1}(z))_{-1}=(\mathrm{proj}_{X_{n}}^{2}(z))_{-1}=a\,\}.$
@@
$\int_{S_{n}^{(a)}}\psi\left(\frac{A_{n}^{b}+B_{n}^{b}}{2}\right)-\frac{1}{2}\psi(A_{n}^{b})-\frac{1}{2}\psi(B_{n}^{b})\,\mathrm{d}\eta_{n}$
$=\int_{{\mathrm{proj}^{-1}_{Z_{n}}}(S_{n}^{(a)})}\psi\left(\frac{A_{n}^{b}\circ\mathrm{proj}_{Z_{n}}+B_{n}^{b}\circ\mathrm{proj}_{Z_{n}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{b}\circ\mathrm{proj}_{Z_{n}})-\frac{1}{2}\psi(B_{n}^{b}\circ\mathrm{proj}_{Z_{n}})\,\mathrm{d}\eta$
$\geq\int_{S^{(a)}}\psi\left(\frac{A_{n}^{b}\circ\mathrm{proj}_{Z_{n}}+B_{n}^{b}\circ\mathrm{proj}_{Z_{n}}}{2}\right)-\frac{1}{2}\psi(A_{n}^{b}\circ\mathrm{proj}_{Z_{n}})-\frac{1}{2}\psi(B_{n}^{b}\circ\mathrm{proj}_{Z_{n}})\,\mathrm{d}\eta,$ (27)
where $A_{n}^{b},B_{n}^{b}$ are as in Theorem 4.6 taking $j_{*}=b$. Observe that
$A_{n}^{b}\circ\mathrm{proj}_{Z_{n}}$ $=\mathbb{E}_{\eta_{n}}\big(\mathbbm{1}_{[b]_{0}}\circ\mathrm{proj}_{X_{n}}^{1}\,\big|\,\mathcal{B}_{1,n}^{-}\vee\mathcal{B}_{0,n}\big)\circ\mathrm{proj}_{Z_{n}}$
$=\mathbb{E}_{\eta}\left(\mathbbm{1}_{[b]_{0}}\circ\mathrm{proj}_{X}^{1}\,\big|\,(\mathrm{proj}_{Z_{n}})^{-1}\left(\mathcal{B}_{1,n}^{-}\vee\mathcal{B}_{0,n}\right)\right)$ (28)
@@
$B_{n}^{b}\circ\mathrm{proj}_{Z_{n}}=\mathbb{E}_{\eta}\left(\mathbbm{1}_{[b]_{0}}\circ\mathrm{proj}_{X}^{2}\,\big|\,(\mathrm{proj}_{Z_{n}})^{-1}\left(\mathcal{B}_{2,n}^{-}\vee\mathcal{B}_{0,n}\right)\right).$ (29)
@@
$\lim_{n\to\infty}\int_{S_{n}^{(a)}}\left[\psi\left(\frac{A_{n}^{b}+B_{n}^{b}}{2}\right)-\frac{1}{2}\psi(A_{n}^{b})-\frac{1}{2}\psi(B_{n}^{b})\right]\mathrm{d}\eta_{n}>0.$ (30)